# DataSculpt: Crafting Data Landscapes for LLM Post-Training through Multi-objective Partitioning

### v1

@@ 5. Conclusion
The ability to process long contexts is crucial for Large Language Models (LLMs), making the effective training of LLMs for extended-context tasks from a data-centric perspective a worthy area of investigation. A pressing challenge lies in organizing long-sequence data for domains lacking extensive text, while ensuring diversity in domain knowledge. In response, we introduce DataSculpt, a framework for organizing long-context training data, conceptualizing the task as a multi-objective combinatorial optimization (MOCO) problem. This framework considers factors including relevance, homogeneity, integrity, and computational costs associated with data construction, employing a coarse-to-fine method for phased greedy solutions. Experimental results indicate that data organized via DataSculpt significantly enhances long-context capabilities without compromising general inference abilities, marking a substantial improvement over existing data management methods.

### v2

@@ 7. Conclusion
The ability to process long contexts is crucial for Large Language Models (LLMs), making the effective training of LLMs for extended-context tasks from a data-centric perspective a worthy area of investigation. A pressing challenge lies in organizing long-sequence data for domains lacking extensive text, while ensuring diversity in domain knowledge. In this work, we introduced DataSculpt, a framework for organizing long-context training data, conceptualizing the task as a multi-objective combinatorial optimization problem. This framework considered factors including relevance, homogeneity, integrity, and computational costs associated with data construction, employing a coarse-to-fine method for phased greedy solutions. Experimental results indicated that data organized via DataSculpt significantly enhances long-context capabilities without compromising general inference abilities, marking a substantial improvement over existing data management methods.